# Slow-Fast Policy Optimization: Reposition-Before-Update for LLM Reasoning

### v2

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Corresponding authors.
@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable progress on complex multi-step reasoning tasks, especially in mathematics problem solving and scientific question answering (OpenAI et al., 2024, DeepSeek-AI et al., 2025b). A central driver of these advances has been reinforcement learning (RL), which fine-tunes LLMs using reward signals tied to semantic correctness and solution quality. For example, DeepSeekMath (Shao et al., 2024) demonstrates that Group Relative Policy Optimization (GRPO) can substantially improve open-source LMs on mathematical reasoning, while DeepSeek-R1 (Guo et al., 2025) shows that RL signals alone can induce emergent reasoning behaviors across various domains.
@@ 2 Preliminaries
Figure 1: Pipeline of SFPO at step $s$. Starting from the current policy $\pi_{\theta^{s,0}}$, we first generate rollouts for training. Stage I (Fast Trajectory): apply $K$ successive gradient updates on the same batch to obtain $\theta^{s,K}$. Stage II (Reposition): interpolate between $\theta^{s,K}$ and the starting point $\theta^{s,0}$ to form $\widetilde{\theta}^{s,K}$, controlling off-policy drift. Stage III (Slow Correction): perform one additional update on $\widetilde{\theta}^{s,K}$, yielding $\pi_{\theta^{s+1,0}}$ for the next step.
@@
Despite its popularity, GRPO inherits structural drawbacks from the underlying on-policy update rule. At each step, the parameters are updated via a single stochastic gradient step estimated from a batch of rollouts. The randomness of rewards leads to high-variance gradient estimates, making updates unstable. Group normalization partially dampens this effect but remains sensitive to fluctuations within each batch. Moreover, restricting each batch to a single update discards potentially useful gradient information across inner steps, leading to inefficient data use and limited variance reduction. These drawbacks underscore the need for an update mechanism that can stabilize gradient directions while making more effective use of available samples.
@@ 3 Method
To address the instability and inefficiency of one-shot policy updates, we propose Slow-Fast Policy Optimization (SFPO), a simple yet general update mechanism that reuses rollouts more effectively while remaining plug-compatible with standard on-policy policy gradient algorithms. Each step consists of three stages: (i) a fast trajectory of multiple inner updates on the same batch, (ii) a reposition step that interpolates back toward the on-policy point to control drift, and (iii) a slow correction via an extra-gradient update. As illustrated in Fig. 1 and Alg. 1, this fast–reposition–slow design transforms noisy one-shot updates into well-structured update trajectories, yielding more stable optimization and higher sample efficiency without additional rollouts.
@@ 3.1 Stage I: Fast Trajectory
In standard on-policy policy-gradient methods such as GRPO, each step is updated by a single stochastic gradient:
@@
Formally, starting from parameters $\theta^{s,0}$ at the beginning of step $s$, we execute a short fast trajectory of $K$ inner updates:
@@ 3.3 Stage III: Slow Correction
  $\alpha\leftarrow\alpha_{0}\cdot\mathbf{1}[\,s<s^{\star}\,]$// Set $\alpha$ for this step from past trigger (non-anticipatory)
@@ 3.4 Scheduling $\alpha$
A nonzero $\alpha$ is essential to exploit the stabilized fast trajectory when $K>0$: if $\alpha=0$, the reposition collapses to $\theta^{s,0}$ and the fast trajectory is discarded, eliminating the benefit of Stage I. However, the same aggressiveness that helps during early training can be counter-productive near a minimizer. When $\|\nabla\mathcal{L}\|$ is large, a nonzero $\alpha$ accelerates progress by moving along the stabilized fast direction; but as we approach a minimum, the signal weakens while curvature and stochastic noise dominate, so a large $\alpha$ amplifies drift and instability. This motivates an adaptive schedule for $\alpha$.
@@
In our implementation, $\alpha$ is scheduled online at each step. After generating rollouts with the current policy, we compute the policy entropy $H_{s}$ and maintain a rolling buffer of the past $\omega$ entropy values (window size $\omega$). Let $\mu_{s}$ and $\sigma_{s}$ denote the mean and standard deviation within this buffer. We define the one-sided z-score
@@
If $|Z_{s}|\geq\tau$ for a threshold $\tau$, we mark the current step $s^{\star}$ and set $\alpha=0$ for all $s\geq s^{\star}$. Otherwise we keep $\alpha=\alpha_{0}$. Intuitively, sharp entropy fluctuations signal that the policy is close to a local optimum where noise dominates, so interpolation should be disabled. This entropy-triggered schedule exploits the fast trajectory during the high-signal early stage of training, while reverting to pure on-policy updates near convergence for stability. [^3] Although a decaying schedule on $\alpha$ could be a formal solution, we find little empirical difference, as shown in Fig. 7 of Sec. 5. Hence, for simplicity and efficiency, we downgrade SFPO to GRPO once the step reaches $s\geq s^{\star}$.
@@ 4.1 Experimental Settings
Reasoning RL Training. We conduct the reasoning RL training on two different training datasets to test the effectiveness of SFPO on different scales of dataset. The first one is the combination of DAPO training dataset (Yu et al., 2025a) and Math training dataset (Hendrycks et al., 2021), which is a total of approximate 24K data. The second one is the Skywork-OR1 Math RL training dataset (He et al., 2025b) with 105k data. The training batch size is set to 256, and the number of responses for each question is set to 8 by default. The total trianing step is set to 400 by default. All the training experiments are done based on verl  (Sheng et al., 2025) with a single 8xH100 GPUs node. More hyperparameters setting for reasoning RL training can be found in  Appendix C.
@@
DS-distilled-Qwen-1.5B Base 73.80 16.67 20.00 47.59 28.86 36.94 37.27
@@
4.2 Main Results
4.2.1 Math Reasoning Benchmarks
As illustrated in Table 1, our proposed SFPO consistently outperforms vanilla GRPO across all base models and benchmarks. Specifically, for small-scale models such as Qwen2.5-Math-1.5B and DS-distilled-Qwen-1.5B, SFPO demonstrates superior performance enhancements on math reasoning benchmarks, raising the average accuracy from 38.35 to 40.19 with an absolute gain of +1.84, and from 47.73 to 50.53 with a gain of +2.80, respectively. The improvements are particularly pronounced on challenging tasks such as AIME24 and AIME25, where DS-distilled-Qwen-1.5B achieves an absolute gain of +7.5 on AIME25. The larger models also exhibit similar performance gains. For Qwen2.5-Math-7B, SFPO raises the average accuracy from 48.36 to 49.19 with an absolute gain of +1.80. For DS-distilled-Qwen-7B, SFPO boosts the average accuracy from 60.47 to 63.04, corresponding to an absolute gain of +0.8. For Qwen3-4B-Base model, SFPO improves average accuracy from 43.99 to 45.59, an absolute gain of +1.60, highlighting its robustness across various models. Moreover, Table 2 demonstrates that SFPO can consistently achieve better training performance compared to GRPO for the larger training dataset Skywork-or1, proving the robustness of SFPO across different scales and distributions of training datasets.
4.2.2 Training Dynamics.
@@ 4.2.1 Math Reasoning Benchmarks
We compare the training dynamics between SFPO and GRPO to better understand the differences in optimization behavior as shown in Fig. 2 and Fig. 3. More comprehensive training dynamics for other models can be found in  Appendix D.
@@
Figure 2: Average validation accuracy of different base models throughout the training process.
Figure 3: Training dynamics for DeepSeek-R1-Distilled-Qwen-7B, comparing GRPO and SFPO across response length, entropy loss, and reward.
@@
Figure 4: Comparison of GRPO and SFPO. (a) Number of rollouts required to achieve the best accuracy of GRPO. (b) Corresponding training time.
We evaluate the efficiency gains of SFPO over GRPO by comparing the total number of rollouts and the wall-clock time required to reach the same benchmark accuracy. As shown in Fig. 4, SFPO consistently surpasses GRPO in rollout efficiency and training speed across all model scales. Specifically, SFPO requires 3.21×, 3.50×, and 4.93× fewer rollouts than GRPO for DS-Qwen-1.5B, Qwen3-4B-Base, and DS-Qwen-7B, respectively, to reach the same best accuracy. This advantage directly translates into reduced training time, where SFPO achieves 2.62×, 2.65×, and 4.19× speedups over GRPO for the same models, significantly lowering the training cost. Note that SFPO does not introduce extra GPU memory overhead as it does not need to store the heavy optimizer status. The detailed profiling results for GPU memory usage in the training process can be found in Appendix B. These significant efficiency gains align with our expectations, since the primary bottleneck in the training process lies in rollout generation, which accounts for more than 70% of the overall inference time (He et al., 2025a). By substantially reducing the number of rollouts required and harnessing the reposition mechanism, SFPO alleviates this bottleneck and achieves faster training.
@@ 6 Related Works
OpenAI O1 (OpenAI, ) introduced a paradigm shift in LLM reasoning by extending the reasoning horizon before final responses. DeepSeek-R1 (Guo et al., 2025) further advanced this line by open-sourcing both its training algorithm, the value-model-free GRPO (Shao et al., 2024), and model weights, achieving performance comparable to O1. Subsequent work has focused on stabilizing and simplifying GRPO: DAPO (Yu et al., 2025b) identifies entropy collapse as a key challenge and proposes effective remedies, while Dr. GRPO (Liu et al., 2025) removes normalization terms without sacrificing performance. In contrast, SFPO is orthogonal to both families: it leaves the objective unchanged yet restructures the update itself into a fast–reposition–slow decomposition, improving variance reduction, stability, and sample efficiency in LLM reasoning.
@@ 7 Conclusion
We proposed Slow-Fast Policy Optimization (SFPO), a simple three-stage update framework that stabilizes early training by combining a fast trajectory, a reposition mechanism, and a slow correction. SFPO directly addresses the instability of one-shot GRPO updates, achieving higher reasoning accuracy and fewer generated tokens without added wall-clock cost. Beyond these gains, its modular structure naturally opens avenues for curriculum or meta-learning extensions, and our entropy-triggered $\alpha$ schedule suggests richer adaptive rules. We believe SFPO offers not only an effective plug-in for current policy optimization, but also a foundation for scalable and data-aware optimization strategies in the LLM reasoning area.

### v3

@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable progress on complex multi-step reasoning tasks, especially in mathematics problem solving and scientific question answering (OpenAI et al., 2024, DeepSeek-AI et al., 2025b). A central driver of these advances has been reinforcement learning (RL), which fine-tunes LLMs using reward signals tied to semantic correctness and solution quality. For example, DeepSeekMath (Shao et al., 2024) demonstrates that Group Relative Policy Optimization (GRPO) can substantially improve open-source LMs on mathematical reasoning, while DeepSeek-R1 (Guo and others, 2025) shows that RL signals alone can induce emergent reasoning behaviors across various domains.
@@ 2 Preliminaries
Figure 1: Pipeline of SFPO at iteration $s$. Starting from the current policy $\pi_{\theta^{s,0}}$, we first generate rollouts for training. Stage I (Fast Trajectory): apply $K$ successive gradient updates on the same batch to obtain $\theta^{s,K}$. Stage II (Reposition): interpolate between $\theta^{s,K}$ and the starting point $\theta^{s,0}$ to form $\widetilde{\theta}^{s,K}$, controlling off-policy drift. Stage III (Slow Correction): perform one additional update on $\widetilde{\theta}^{s,K}$, yielding $\pi_{\theta^{s+1,0}}$ for the next iteration.
@@
Despite its popularity, GRPO inherits structural drawbacks from the underlying on-policy update rule. At each iteration, the parameters are updated via a single stochastic gradient step estimated from a batch of rollouts. The randomness of rewards leads to high-variance gradient estimates, making updates unstable. Group normalization partially dampens this effect but remains sensitive to fluctuations within each batch. Moreover, restricting each batch to a single update discards potentially useful gradient information across inner steps, leading to inefficient data use and limited variance reduction. These drawbacks underscore the need for an update mechanism that can stabilize gradient directions while making more effective use of available samples.
@@ 3 Method
To address the instability and inefficiency of one-shot policy updates, we propose Slow-Fast Policy Optimization (SFPO), a simple yet general update mechanism that reuses rollouts more effectively while remaining plug-compatible with standard on-policy policy gradient algorithms. Each iteration consists of three stages: (i) a fast trajectory of multiple inner updates on the same batch, (ii) a reposition step that interpolates back toward the on-policy point to control drift, and (iii) a slow correction via an extra-gradient update. As illustrated in Fig. 1 and Alg. 1, this fast–reposition–slow design transforms noisy one-shot updates into well-structured update trajectories, yielding more stable optimization and higher sample efficiency without additional rollouts.
@@ 3.1 Stage I: Fast Trajectory
In standard on-policy policy-gradient methods such as GRPO, each outer iteration is updated by a single stochastic gradient:
@@
Formally, starting from parameters $\theta^{s,0}$ at the beginning of iteration $s$, we execute a short fast trajectory of $K$ inner updates:
@@ 3.3 Stage III: Slow Correction
  $\alpha\leftarrow\alpha_{0}\cdot\mathbf{1}[\,s<s^{\star}\,]$// Set $\alpha$ for this iteration from past trigger (non-anticipatory)
@@
  Compute entropy $H_{s}$; update rolling buffer $\mathcal{H}$ (keep last $\omega$ ones) and $(\mu_{s},\sigma_{s})$.
@@ 3.4 Scheduling $\alpha$
A nonzero $\alpha$ is essential to exploit the stabilized fast trajectory when $K>0$: if $\alpha=0$, the reposition collapses to $\theta^{s,0}$ and the fast trajectory is discarded, eliminating the benefit of Stage I. However, the same aggressiveness that helps early can be counter-productive near a minimizer. When $\|\nabla\mathcal{L}\|$ is large, a nonzero $\alpha$ accelerates progress by moving along the stabilized fast direction; but as we approach a minimum, the signal weakens while curvature and stochastic noise dominate, so a large $\alpha$ amplifies drift and instability. This motivates an adaptive schedule for $\alpha$.
@@
In our implementation, $\alpha$ is scheduled online at each iteration. After generating rollouts with the current policy, we compute the policy entropy $H_{s}$ and maintain a rolling buffer of the past $\omega$ entropy values (window size $\omega$). Let $\mu_{s}$ and $\sigma_{s}$ denote the mean and standard deviation within this buffer. We define the one-sided z-score
@@
If $|Z_{s}|\geq\tau$ for a threshold $\tau$, we mark the current step $s^{\star}$ and set $\alpha=0$ for all $s\geq s^{\star}$. Otherwise we keep $\alpha=\alpha_{0}$. Intuitively, sharp entropy fluctuations signal that the policy is close to a local optimum where noise dominates, so interpolation should be disabled. This entropy-triggered schedule exploits the fast trajectory in the high-signal early phase, while reverting to pure on-policy updates near convergence for stability. [^3] Although a decaying schedule on $\alpha$ could be a formal solution, we find little empirical difference, as shown in Fig. 7 of Sec. 5. Hence, for simplicity and efficiency, we downgrade SFPO to GRPO once the iteration reaches $s\geq s^{\star}$.
@@ 4.1 Experimental Settings
Reasoning RL Training. We conduct the reasoning RL training on two different training datasets to test the effectiveness of SFPO on different scales of dataset. The first one is the combination of DAPO training dataset (Yu et al., 2025a) and Math training dataset (Hendrycks et al., 2021), which is a total of approximate 24K data. The second one is the Skywork-OR1 Math RL training dataset (He et al., 2025b) with 105k data. The training batch size is set to 256, and the number of responses for each question is set to 8 by default. The total trianing step is set to 400 by default. All the training experiments are done based on verl  (Sheng et al., 2025) with a single 8-GPU node.
@@
4.2 Main Results
4.2.1 Math Reasoning Benchmarks
As illustrated in Table 1, our proposed SFPO consistently outperforms vanilla GRPO across all base models and benchmarks. Specifically, for small-scale models such as Qwen2.5-Math-1.5B and DS-distilled-Qwen-1.5B, SFPO demonstrates superior performance enhancements on math reasoning benchmarks, raising the average accuracy from 38.35 to 40.19 with an absolute gain of +1.84, and from 47.73 to 50.53 with a gain of +2.80, respectively. The improvements are particularly pronounced on challenging tasks such as AIME24 and AIME25, where DS-distilled-Qwen-1.5B achieves an absolute gain of +7.5 on AIME25. The larger models also exhibit similar performance gains. For Qwen2.5-Math-7B, SFPO raises the average accuracy from 48.36 to 49.19 with an absolute gain of +0.83. For DS-distilled-Qwen-7B, SFPO boosts the average accuracy from 60.47 to 63.04, corresponding to an absolute gain of +2.57. For Qwen3-4B-Base model, SFPO improves average accuracy from 43.99 to 45.59, an absolute gain of +1.60, highlighting its robustness across various models. Moreover, Table 2 demonstrates that SFPO can consistently achieve better training performance compared to GRPO for the larger training dataset Skywork-or1, proving the robustness of SFPO across different scales and distributions of training datasets.
@@ 4.2.1 Math Reasoning Benchmarks
DS-distilled-Qwen-1.5B Base 73.80 16.67 20.00 47.59 28.86 36.94 37.31
@@
4.2.2 Training Dynamics.
We compare the training dynamics between SFPO and GRPO to better understand the differences in optimization behavior as shown in Fig. 2 and Fig. 3.
@@
Figure 2: Average validation accuracy of different base models throughout the learning process.
Figure 3: Comparison of training behaviors in terms of response length, entropy loss, and reward.
Figure 4: Comparison between GRPO and SFPO. (a) Number of rollouts required to achieve the same best accuracy as GRPO. (b) Corresponding training time.
@@
We evaluate the efficiency gains of SFPO over GRPO by comparing the total number of rollouts and the wall-clock time required to reach the same benchmark accuracy. The results in Fig. 4 illustrate that SFPO consistently outperforms GRPO in both rollout efficiency and training time across all model scales. To be specific, SFPO requires 3.21×, 3.50×, and 4.93× fewer rollouts than GRPO for DS-Qwen-1.5B, Qwen3-4B-Base, and DS-Qwen-7B, respectively, to reach the same best accuracy. This advantage directly translates into reduced training time, where SFPO achieves 2.62×, 2.65×, and 4.19× speedups over GRPO for the same models, significantly lowering the training cost. Note that SFPO does not introduce extra GPU memory overhead as it does not need to store the heavy optimizer status. The detailed profiling results for GPU memory usage in the training process can be found in Appendix B. These significant efficiency gains align with our expectations, since the primary bottleneck in the training process lies in rollout generation, which accounts for more than 70% of the overall inference time (He et al., 2025a). By substantially reducing the number of rollouts required and harnessing the reposition mechanism, SFPO alleviates this bottleneck and achieves faster training.
@@ 6 Related Works
OpenAI O1 (OpenAI, ) introduced a paradigm shift in LLM reasoning by extending the reasoning horizon before final responses. DeepSeek-R1 (Guo and others, 2025) further advanced this line by open-sourcing both its training algorithm, the value-model-free GRPO (Shao et al., 2024), and model weights, achieving performance comparable to O1. Subsequent work has focused on stabilizing and simplifying GRPO: DAPO (Yu et al., 2025b) identifies entropy collapse as a key challenge and proposes effective remedies, while Dr. GRPO (Liu et al., 2025) removes normalization terms without sacrificing performance. In contrast, SFPO is orthogonal to both families: it leaves the objective unchanged yet restructures the update itself into a fast–reposition–slow decomposition, improving variance reduction, stability, and sample efficiency in LLM reasoning.
@@ 7 Conclusion
We proposed Slow-Fast Policy Optimization (SFPO), a simple three-stage update mechanism that stabilizes early training by combining a fast trajectory, a reposition step, and a slow correction. SFPO directly addresses the instability of one-shot GRPO updates, achieving higher reasoning accuracy and fewer generated tokens without added wall-clock cost. Beyond these gains, its modular structure naturally opens avenues for curriculum or meta-learning extensions, and our entropy-triggered $\alpha$ schedule suggests richer adaptive rules. We believe SFPO offers not only an effective plug-in for current policy optimization, but also a foundation for scalable and data-aware optimization strategies in the LLM reasoning area.
Acknowledgments
Huo is partially sponsored by a subcontract of NSF grant 2229876, the A. Russell Chandler III Professorship at Georgia Institute of Technology, an NIH-sponsored Georgia Clinical & Translational Science Alliance, and the Georgia Department of Transportation.
Reproducibility Statement
All results reported in Sec. 4 are fully reproducible. We will release the code and experiment scripts upon acceptance.